determine how many liters of a 7 percent acid solution and how many liters of a 12 percent acid solution should be mixed by a chemist to obtain 6 liters of a 10 percent acid solution.

Answer by chessace(471)

x * 0.07 + y * 0.12 = 6 * 0.10
x + y = 6
7x + 12y = 60
7x + 7y = 42
5y = 18
y = 3.6
x = 2.4